OpenStudy (anonymous):

Prove that the function y =(x^4 - x^3 + x^2 - x -2)/(x^2 + x + 1) crosses the x-axis somewhere between -1 and 1

OpenStudy (anonymous):

and the unit we are workin on is derivatives

OpenStudy (anonymous):

when x=1, y=-2/3 when x=-1, y=2. Therefore, there must be y=0 between -1<=x<=1

OpenStudy (anonymous):

but that might be a discontinuity

OpenStudy (anonymous):

I believe you want to use the intermediate value theorem. When you plug in x = -1, you get a positive answer, and when you plug in x = 1 you get a negative answer. Because the function is continuous between -1 and 1, you can use the intermediate value theorem to say there must be some point inbetween where the value of the function is 0.

OpenStudy (anonymous):

I think because they are polynomials they are continuous for all x

OpenStudy (anonymous):

how do you know it's continuous there

OpenStudy (anonymous):

x^2+x+1 = (x+1/2)^2 + 3/4, and is always larger than 3/4

OpenStudy (anonymous):

there are no points of discontinuity in the reals. the values of x that make x^2+x+1 =0 are complex (cube roots of unity to be exact)

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

have you done the maple part of the assignment yet?

OpenStudy (anonymous):

im on part b) it wont work for me tho

OpenStudy (anonymous):

what did you graph?

OpenStudy (anonymous):

f(x)= x^2-1-sinx

OpenStudy (anonymous):

yeah that's what I did too

OpenStudy (anonymous):

just copy and pasted what they did in the pre lab for part b but it didnt work

OpenStudy (anonymous):

that's what I did too, even the interval for the first one was the same. It worked for me

OpenStudy (anonymous):

it wouldnt solve for f(a) and f(d)

OpenStudy (anonymous):

I used the loop to solve that part. Do you think we're supposed to do all 10 iterations?

OpenStudy (anonymous):

no i just copy and pasted the loop part

OpenStudy (anonymous):

did you assign an a and a b variable before running the loop?

OpenStudy (anonymous):

ya then (a+b)/2=d

OpenStudy (anonymous):

you don't need to do a d since that is already in the loop

OpenStudy (anonymous):

try going through each of your expression statements and pressing enter on all of them again. Sometimes that works when it is messing up. or copy and paste the loop in again

OpenStudy (anonymous):

k ya i tried everything- how is the rest of your assignment?

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

number 4 i was havin trouble with- the other ones look like they are pretty much from our notes tho

OpenStudy (anonymous):

I've actually just started lol. I just finished number one when I read your question.

OpenStudy (anonymous):

ya its not a polynomial so i dunno how to prove its continuous

OpenStudy (anonymous):

for#1

OpenStudy (anonymous):

Well the bottom will never be 0 because of the +1 right

OpenStudy (anonymous):

just could be factored

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

you still need to keep in mind the original restrictions on the function. The graph of x^2+x+1 is always positive. So the function would be defined anywhere.

OpenStudy (anonymous):

ok so its simple then-on to #2

OpenStudy (anonymous):

I haven't even really started that one, but I never did linearization in class

OpenStudy (anonymous):

i know eh - they really need to stop screwin with us and stick to what they taught us

OpenStudy (anonymous):

at least the last one was pretty easy, and from what you say this one is pretty straight forwards to right?

OpenStudy (anonymous):

the one before that was really hard though and there are always those problems that you've never seen before too

OpenStudy (anonymous):

ya man number 2 and 3 are pretty easy

OpenStudy (anonymous):

what about 5 and 6?

OpenStudy (anonymous):

5 looks hard but 6 looks like its from our notes

OpenStudy (anonymous):

actually i dont know how to do 3

OpenStudy (anonymous):

for a you take the lim as x-> +inf and whatever that is is the asymptote right? for b we have that equation of the tangent line formula for c you can use a online graphing thing to get a better idea of what it should look like I think, just from a glance through anyways. I'll get there after number 2 though. Maybe I'll just do it tomorrow lol

OpenStudy (anonymous):

ya me too haha- last resort ill just come on here and post everything

OpenStudy (anonymous):

sometimes I've come on here and noticed you already asked the question I was going to ask and have already got an answer lol

OpenStudy (anonymous):

ya man just check mine first- i dont think its cheating cuz im learning too

OpenStudy (anonymous):

I think it's the same as if we went and asked a TA. People on here are just more convenient to get help from

OpenStudy (anonymous):

k im done for tonite

OpenStudy (anonymous):

same lol, night!

OpenStudy (anonymous):

night